silvernight269:

Write the equation of the polynomial of lowest possible degree in factored form. Has integer coefficients, one zero at x=1/2-5i and another at x=-2-√3

jhonyy9:

so i think this ,,one zero " mean root so x_1 and x_2 and we know that a quadratic equation has 2 roots using the Viéte formulas 1. x_1+x_2=−b/a ​ and x_1⋅x_2=c/a 2. a⋅(x-x_1)(x-x_2)=0. 3. ax2+bx+c=a⋅(x-x_1)⋅(x-x_2)=0. hope helped understandably